If in two circles, arcs of same length, subtend angles 120o and 150o at the centre, find the ratio of their radii.

Grade:

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
we know that l=r\Theta
as arcs have same length
r1/r2=\Theta2/\Theta1
so r1/r2=5/4.
Thanks and regards
Shaik Aasif
askIITians faculty